OpenStudy (anonymous):

Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2).

OpenStudy (anonymous):

Perpendicular lines of a slope that, if multiplied with the slope of a line, equals -1. That is: (slope of line) * (slope of perpendicular line) = -1 So, if you know the slope of the line you start with, you can find the slope of the perpendicular line with this. Then you can use the point-slope form of a line to find the final equation.

OpenStudy (anonymous):

So, I have to get the X put of my options, and whichever equals -1 is my answer?

OpenStudy (anonymous):

First, find the slope of the line in the graph.

OpenStudy (anonymous):

(-2, 0) (0, -1) \[Y = mx+b\] Don't I put my X values in?

OpenStudy (anonymous):

\[slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

Ohhh ,whoops. \[\frac{ ^{-}1 - 0 }{ 0 - ^{-}2 } = -\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

There ya go. Now you can find the slope of the line that is perpendicular to that. (slope of line)*(slope of perpendicular line) = -1

OpenStudy (anonymous):

\[-\frac{ 1 }{ 2 } \div -1 = \] slope of perpendicular line?

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

let "a" be the slope of the line we know, and "m" be the slope of the perpendicular line: \[a*m = -1\] \[\frac{-1}{2} * m = -1\] \[m = -1 * \frac{2}{-1} = 2\]

OpenStudy (anonymous):

Okay, so \[Y = -\frac{ 1 }{ 2 }x + 2?\]

OpenStudy (anonymous):

No. We just found the SLOPE of the perpendicular line. Now we need to use the point-slope form of a line to find the final equation. \[y - y_o = m(x - x_o)\]

OpenStudy (anonymous):

What are the subscripts?

OpenStudy (anonymous):

They just tell us that they are values for a point. Those are where you'll plug in your given x and y (the point that the line passes through)

OpenStudy (anonymous):

Am I replacing \[y - y _{o}\] and \[x - x_o\] with my X and Y values?..

OpenStudy (anonymous):

Like (-2, 0) and (0, -1)?

OpenStudy (anonymous):

The point you are given for the line to go through is in the form: \[(x_o , y_o) \] You'll only replace the x and y that have the subscript, and then solve for the y that doesn't.

OpenStudy (anonymous):

You are given the point (-1,2), and you have a slope of 2. \[y - 2 = 2(x - (-1))\] Solve for y, and you're done.

OpenStudy (anonymous):

\[y = 2x + 4\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

Thank you!!